OpenStudy (kanwal32):

how to break 1/(2n+1)(2n+3)(2n+5) into partial fraction

OpenStudy (kanwal32):

\[\frac{ 1 }{ (2n+1)(2n+3)(2n+5) }\]

OpenStudy (dls):

Let.. \[\Large \frac{ 1}{(2n+1)(2n+3)(2n+5) } = \frac{A}{2n+1} + \frac{B}{2n+3} + \frac{C}{2n+5}\]

OpenStudy (dls):

take LCM..solve for A B and C..substitute

OpenStudy (kanwal32):

ok, got it :)

OpenStudy (dls):

glad :)

OpenStudy (kanwal32):

can i solve it by multiplying by (2n+1)

OpenStudy (kanwal32):

@hartnn

hartnn (hartnn):

what do u mean ? why only multiply by 2n+1 ?

OpenStudy (kanwal32):



OpenStudy (kanwal32):

now substituting by n by -1/2 so getting the value of A

hartnn (hartnn):

but that won't help you get A,B,C values

OpenStudy (kanwal32):

repeating the

OpenStudy (kanwal32):

process again

hartnn (hartnn):

ok, you can do that

hartnn (hartnn):

after taking LCM also, you'd put x=-1/2 to find A, its same thing, just different ways of doing.